You have an outstanding credit card balance of $2,000. You decide to stop making new charges and pay $375 each month until the balance is paid. If your annual interest rate is 18%, how many months will it take you to pay off your balance?
Answer #1

PV =

[1-(4+5)-) [1-(1+r)-n LT P= Periodic Payment r=rate per period n = number of periods

r = 18%/12 = 1.5% (per month)

2000 = 375 41-(1 + 0.015)-n 0.015

0.08 = 1 - (1.015)-n

0.92 =(1.015)-n

Taking log of both sides

LN(0.92) = -n * LN(1.015)

-0.083382 = - n * 0.014889

n = 5.60 months